A ball is thrown from a cliff the cliff is 70 m high the ball reaches a maximum height of 100 m it reaches the ground at a .90 m from where it started
1
Expert's answer
2020-09-15T17:23:58-0400

Let us write the equations of motion:

"x = v_0\cos\alpha t, \;\; y = y_0 + v_0\sin\alpha t - \dfrac{gt^2}{2}."

So, "t = \dfrac{x}{v_0\cos\alpha}, \;\; y = y_0 + x\tan\alpha - \dfrac{g}{2v_0^2\cos^2\alpha}x^2" .

The last equation is an equation of a parabola.

When x = 90, y = 0, when x = 0, y = 70, and the ordinate of a vertex is 100.

So, "y = 70 + x\tan\alpha - \dfrac{5}{v_0^2\cos^2\alpha}x^2" .

"0 = 70 +90\tan\alpha - \dfrac{5}{v_0^2\cos^2\alpha}\cdot90^2" (1)

The abscissa of a vertex is "x_0 = -\dfrac{\tan\alpha}{-2\frac{5}{v_0^2\cos^2\alpha}} = \dfrac{v_0^2\cos^2\alpha\tan\alpha}{10}." And "100 = 70 + \dfrac{v_0^2\cos^2\alpha\tan\alpha}{10}\tan\alpha - \dfrac{v_0^2\cos^2\alpha\tan^2\alpha}{20} = 70 + \dfrac{v_0^2\cos^2\alpha\tan^2\alpha}{20} ." (2)


Therefore, (1) "0 = 7 +9\tan\alpha - \dfrac{5}{v_0^2\cos^2\alpha}\cdot810."

(2) "30 = \dfrac{v_0^2\cos^2\alpha\tan^2\alpha}{20} \Rightarrow v_0^2 = \dfrac{600}{\cos^2\alpha\tan^2\alpha}"


(1) "0 = 7 +9\tan\alpha - \dfrac{5\cos^2\alpha\tan^2\alpha}{600\cos^2\alpha}\cdot810, \;\;
0 = 7+9\tan\alpha - \dfrac{27}{4}\tan^2\alpha."

"\tan\alpha \approx 1.88, \alpha \approx 62^\circ."

"v_0\approx" 27.7 m/s.